# Memory-adaptive Depth-wise Heterogenous Federated Learning

### v2

@@ Introduction
where ${\cal L}$ is a loss function, e.g, cross-entropy; $\{z_{j-1,i}^{t+1}\}_{i=1}^{n_{k}}$ are activations obtained after the local training samples forward-passed through the first $j$ blocks, i.e., $z_{j-1,i}^{t+1}=f(x_{i};\{\theta_{k,\ell}^{t+1}\}_{\ell=1}^{j-1})$ for $i\in[n_{k}]$ and $f(\cdot;\{\theta_{k,\ell}^{t+1}\}_{\ell=1}^{j-1})$ is the neural network up to the first $j-1$ blocks; $n_{k}$ is the total number of training samples. Specifically, $z_{0,i}^{t+1}=x_{i}$ for all $i\in[n_{k}]$. Problem (1) can be solved by any stochastic gradient-based method. When solving for the $j$-th subproblem at the $t$th round, to fully leverage the global model ${\cal W}^{t}$’s body and the locally newly updated head, we use $(\theta_{j}^{t},\phi_{j-1}^{t+1})$ as the initial point. See the data flow in Figure 4 when performing the training for the $j$th block. We remark that when the memory budget permits, the activation $\{z_{j-1,i}^{t+1}\}_{i=1}^{n_{k}}$, which no longer requires the gradient, could be buffered to compute $\{z_{j,i}^{t+1}\}_{i=1}^{n_{k}}$ once the $\theta_{k,j}^{t+1}$ is obtained, hence saving the redundant forward-pass from the first block to the $j$th block. Also, the number computation required for approximately solving $J_{k}$ subproblems in the form of (1) should be similar to approximately solving the problem $\min_{{\cal W}}F_{k}({\cal W})$ if we perform the same number of local updates, and ignore the negligible computation overhead in updating in the head $\phi$. This is because the amount of computation required by one gradient evaluation $\nabla_{{\cal W}}F_{k}$ is equivalent to that of the summation of gradient evaluation of $\{\nabla_{\theta_{j}}{\cal L}\}_{j=1}^{J_{k}}$.
@@
[…]
@@
[…]
@@
Table 2: Test results (top-1 accuracy) under balanced non-IID data partitions using PreResNet-20. Grey texts indicate that the training cannot conform to the pre-defined budget constraint. If not specified, FedAvg denotes the results with ×$\times\min(r)$width network. We highlight the best results with Blue Shadow, Red Shadow, and Bold in the scenarios including clients equipped with fairly sufficient, insufficient and abundant memory, respectively.
@@
Unrealistic FedAvg (×$\times 1$) 63.44 $\pm$ 2.95 69.77 $\pm$ 1.48 34.10 $\pm$ 1.68 68.33 $\pm$ 0.75 28.14 $\pm$ 0.48 31.28 $\pm$ 0.17 31.56 $\pm$ 0.28 43.54 $\pm$ 0.16
Fair FedAvg (×$\times\frac{1}{6}$) 47.66 $\pm$ 2.67 53.10 $\pm$ 1.27 25.82 $\pm$ 1.34 54.02 $\pm$ 0.96 15.83 $\pm$ 0.25 16.39 $\pm$ 0.08 13.89 $\pm$ 0.21 17.96 $\pm$ 0.08
@@
Table 2: Test results (top-1 accuracy) under balanced non-IID data partitions using PreResNet-20. Grey texts indicate that the training cannot conform to the pre-defined budget constraint. If not specified, FedAvg denotes the results with ×$\times\min(r)$width network. We highlight the best results with Blue Shadow, Red Shadow, and Bold in the scenarios including clients equipped with fairly sufficient, insufficient and abundant memory, respectively.

### v3

# Memory-adaptive Depth-wise Heterogeneous Federated Learning

@@ Introduction
[…]
@@
B1∼3 20.02 $\times\frac{1}{8}$ 14.51
B4 14.05 $\times\frac{1}{6}$ 19.34
B5∼6 10.07 $\times\frac{1}{3}$ 38.68
B7 7.21 $\times\frac{1}{2}$ 58.02
B8∼9 5.28 $\times 1$ 116.04
Table 1: Memory cost (in MB) with respect to depth and width of PreResNet-20. Each block consists of 2 convolution layers. B1∼3 indicates Block $1,2$ and $3$ in PreResNet-20 have the same memory cost of 20.02 MB. The values are estimated by pytorch-summary [^1] https://github.com/sksq96/pytorch-summary.
@@
Memory budgets. Using Pre-Activation ResNet-20 (PreResNet-20) (He et al. 2016b) as an example, we show the relation of the memory cost of each block between width-wise and depth-wise training in Table 1. We can see that if clients afford to train $\frac{1}{6}$-width PreResNet-20, they can train the full-size neural network via depth-wise training. The training order is { $_{1}\rightarrow B$$_{2}\rightarrow$ $_{3}\rightarrow B$$_{4}\rightarrow$ $_{5,6}\rightarrow B$7,8,9 }. Inspired by this example, we simulate three memory budget scenarios.
@@
Unrealistic FedAvg ($\times 1$) 63.44 $\pm$ 2.95 69.77 $\pm$ 1.48 34.10 $\pm$ 1.68 68.33 $\pm$ 0.75 28.14 $\pm$ 0.48 31.28 $\pm$ 0.17 31.56 $\pm$ 0.28 43.54 $\pm$ 0.16
Fair FedAvg ($\times\frac{1}{6}$) 47.66 $\pm$ 2.67 53.10 $\pm$ 1.27 25.82 $\pm$ 1.34 54.02 $\pm$ 0.96 15.83 $\pm$ 0.25 16.39 $\pm$ 0.08 13.89 $\pm$ 0.21 17.96 $\pm$ 0.08
@@
Table 2: Test results (top-1 accuracy) under balanced non-IID data partitions using PreResNet-20. Grey texts indicate that the training cannot conform to the pre-defined budget constraint. If not specified, FedAvg denotes the results with $\times\min(r)$ -width network. We highlight the best results with Blue Shadow, Red Shadow, and Bold in the scenarios including clients equipped with fairly sufficient, insufficient and abundant memory, respectively.
@@
Global Model Evaluation
Results in the Fair Budget scenario. In Table 2, we compare test results on a global test dataset (10,000 samples) considering a variety of balanced non-IID data partition and memory constraints. We highlight the best results under different scenarios. In all cases, HeteroFL, SplitMix, and FeDepth family outperform vanilla FedAvg, showing their system designs’ effectiveness under balanced data distribution. Among all methods, our proposed FeDepth and $m$-FeDepth achieve the best performance with significant improvements. For example, on CIFAR10, under Fair Budget, FeDepth gains $4.09\pm 0.30\%$ average improvement compared to HeteroFL while gains $3.99\pm 1.77\%$ average improvement compared to SplitMix. $m$-FeDepth gains $5.35\pm 1.13\%$ average improvement compared to HeteroFL while gains $5.25\pm 1.20\%$ average improvement compared to SplitMix. Figure 6 shows convergence curves of FeDepth on non-IID CIFAR-10 dataset.
Figure 6: Convergence of FeDepth family on Cifar10.
Results in the Lack Budget scenario. We observe that HeteroFL has relatively slight accuracy drops or increases compared to the fair budget scenario. The explanation could be deducted from the behaviors of sub-networks discussed in the previous section and Figure 2 that small sub-networks slightly influence the global performance because the small number of model parameters provides limited knowledge of the global model in the aggregation phase of FL. In contrast, SplitMix has an apparent performance degradation of an average of $5.55\pm 0.81\%$ due to the weaker base learner. The FeDepth and $m$-FeDepth are relatively stable algorithms against insufficient memory budget, showing $1.73\pm 0.34\%$ and $2.74\pm 0.97\%$ degradation, respectively.
Results in the Surplus Budget scenario. We let the new clients with rich resources $r=2$ join in FL and replace the clients with $r=1$. Prior works, like HeteroFL and SplitMix, did not consider such dynamics, and the clients with more memory budgets still train $\times 1$ neural networks. An alternative way is to train a new large base model from scratch and discard previously trained $\times 1$ neural networks, hence wasting computing resources.
From Table 2, we can observe that MKD indeed makes sense for improving the performance of the global model (still $\times 1$-width). Furthermore, we note that combining depth-wise training and MKD is a flexible solution to simultaneously solve dynamic partition, device upgrade, and memory constraints. For example, when a new client with $r=\frac{7}{6}$ enters into the federated learning, the client can locally learn two models via regular and depth-wise sequential training, respectively, and then perform MKD while maintaining an original-size model for aggregation.
Comparison between FeDepth and its variant. As shown in Table 2, for CIFAR-10, FeDepth and $m$-FeDepth can achieve similar prediction accuracy. However, for CIFAR-100, $m$-FeDepth always outperforms FeDepth. It is worth recalling the design of FeDepth, which introduces zero paddings to match the dimension between two skip-connected blocks. This may inject negligible noise for the training on more complex data. Replacing the zero paddings with other modules, such as convolutions, may result in a better model. However, this usually comes at the cost of extra memory because of the new activations and parameters, which is usually intolerable to resource-constrained devices.
Figure 7: Fine-tuning ViT-T/16 on CIFAR-10 (left) and CIFAR-100 (right) under balanced non-IID data partitions with FedAvg FeDepth, and $m$-FeDepth. FedAvg $(\times 1)$ assumes each client can afford to train the full-size model with 12 identical encoder blocks, while FedAvg $(\times\frac{1}{6})$ assumes each client trains a $\frac{1}{6}$-width model, whose memory consumption is equal to train two encoder blocks.
Influence of Unbalanced Non-IID Distribution
Table 3 shows the prediction results on distributed datasets in FL from the unbalanced Dirichlet partition (Fair Budget). We note that HeteroFL and SplitMix were not evaluated on such an unbalanced distribution. Overall, the higher skewed distribution leads to worse performance for FL, which can be observed by comparing results on Table 2 and Table 3.
Since the number of samples per class in CIFAR-100 is limited (there are 500 samples for each class), $\alpha_{u}(\lambda)$ and $\alpha(\lambda)$ will output similar statistical distribution according to the number of samples on each client in FL. Therefore, we obtain similar CIFAR-100 results on both balanced and unbalanced non-IID data partitions. Specifically, $\alpha(\lambda)$ always outputs $400$ training samples per client on average. For CIFAR-100, $\alpha_{u}(0.3)$ outputs $399.40\pm 34.53$ training samples per client, $\alpha_{u}(1.0)$ outputs $399.34\pm 17.74$. For CIFAR-10, $\alpha_{u}(0.3)$ outputs $399.44\pm 150.60$ training samples per client, $\alpha_{u}(1.0)$ outputs $399.39\pm 77.37$.
Regarding CIFAR-10 results, we observe that HeteroFL and SplitMix cannot achieve comparable predictions or generalization ability compared to FedAvg. SplitMix even performs worse than training with the smallest models in FL. This result indicates that SplitMix is not robust to unbalanced distribution. One reason for this phenomenon is that small base models cannot capture representative features due to the significant weight divergence between local clients stemming from a highly skewed distribution (Frankle and Carbin 2019; Li et al. 2022). For HeteroFL, as mentioned in the case study in Section Behaviors of Sub-networks in Prior Works, the full-size neural networks on resource-sufficient devices provide the fundamental ability but small sub-networks trained with unbalanced distribution indeed affect the global performance. In contrast to HeteroFL and SplitMix, our proposed FeDepth and $m$-FeDepth gain substantial improvements of $6.15\%$ and $6.53\%$ on CIFAR-10, and of $12.24\%$ and $11.57\%$ on CIFAR-100 compared to FedAvg.
CIFAR-10 CIFAR-100
Method $\alpha_{u}(0.3)$ $\alpha_{u}(1.0)$ $\alpha_{u}(0.3)$ $\alpha_{u}(1.0)$
FedAvg ($\times\frac{1}{6}$) 46.46 52.02 15.62 17.99
HeteroFL 46.14 52.20 16.02 18.36
SplitMix 31.23 44.70 22.68 25.28
DepthFL 47.13 55.49 23.19 26.02
FeDepth 52.61 58.55 23.25 26.16
$m$-FeDepth 51.58 57.91 27.86 29.56
Table 3: Experimental results on unbalanced Dirichlet partitions. Because of the relatively limited number of samples per class in CIFAR-100, an unbalanced Dirichlet partition outputs similar statistical distribution according to the number of local data examples.
Depth-wise Fine-tuning on ViT
Foundation models or Transformer architectures (Vaswani et al. 2017; Zhou et al. 2023), such as Vision Transformer (ViT) (Dosovitskiy et al. 2020), has shown robustness to distribution shifts (Bhojanapalli et al. 2021). Recent work has demonstrated that replacing a convolutional network with a pre-trained ViT can greatly accelerate convergence and result in better global models in FL(Qu et al. 2022). Inspired by this finding, we hypothesize that fine-tuning ViT with depth-wise learning will still produce a better global model because 1) decomposing blocks in a depth-wise manner maintains the knowledge learned from pretraining, and 2) memory consumptions of activations in each ViT’s block are identical, which indicates that skip connection for handling resource constraints does not introduce any noises and extra parameters. The memory budgets in terms of the width shrinkage ratio $r=\{\frac{1}{6},\frac{1}{3},\frac{1}{2},1\}$ are uniformly allocated to 100 clients as the same setting of PreResNet-20 in the scenario of Fair Budget.
For fine-tuning, we choose a learning rate of $5\times 10^{4}$ and a training epoch of 100. Figure 7 shows the test results of ViT-T/16 (Qu et al. 2022) under balanced Dirichlet data partitions, on which we observe that exploiting FeDepth and $m$-FeDepth can produce good global models. Specifically, FeDepth-ViT significantly outperforms FeDepth-PreResNet-20 with $36.06\pm 12.79\%$ and $30.64\pm 4.24\%$ improvements on CIFAR-10 and CIFAR-100 on average, respectively. $m$-FeDepth-ViT significantly outperforms $m$-FeDepth-PreResNet-20 with $36.66\pm 12.88\%$ and $27.41\pm 3.13\%$ improvements on CIFAR-10 and CIFAR-100 on average, respectively. We also observe that although local ViTs are fine-tuned on varying distribution data, we obtain global models with similar performance. It indicates that ViT is more robust to distribution shifts and hence improves FL over heterogeneous data.
Conclusions
Despite the recent progress in FL, memory heterogeneity still remains largely underexplored. Unlike previous methods based on width-scaling strategies or fixed-depth split, we propose adaptive depth-wise learning for handling varying memory capabilities. The experimental results demonstrate our proposed FeDepth family outperform the state-of-the-art algorithms including HeteroFL, SplitMix and DepthFL and are robust to data heterogeneity and client participation. Furthermore, using the robustness of ViT to heterogeneous distribution shifts, we reach an excellent global model via our depth-wise solutions. FeDepth is a flexible and scalable framework that can be compatible with most FL algorithms and is reliable to be deployed in practical FL systems and applications.
Appendix A Visualization of Label Distribution
We consider 100 clients in all experiments, and in Figure 8, we show label distributions of 5 out of 100 clients under balanced $\alpha(0.3),\alpha(1.0)$ and unbalanced $\alpha_{u}(0.3),\alpha_{u}(1.0)$ splits of CIFAR-10.
Appendix B Extensive Experimental Results
Large-scale FL experiments
We also conducted experiments on EMNIST with 500 and 1000 clients, respectively, with the 0.1 participation rate, and fair budget with $\alpha(1)$. Additionally, we report the results on FEMNIST (Caldas et al. 2018), a natural-split FL dataset derived from partitioning 3597 writers from EMNIST. Furthermore, we present results on TinyImageNet under with 100 clients and 0.1 participation rate. The results are shown in the following table.
Datasets FedAvg HeteroFL SplitMix DepthFL FeDepth $m$-FeDepth
EMNIST (0.5K) 77.44 79.26 70.94 73.88 82.07 81.73
EMNIST (1.0K) 74.24 77.95 62.04 70.18 81.91 81.68
FEMNIST 62.69 71.05 54.62 73.80 78.07 76.24
TinyImageNet 21.00 23.98 30.87 30.89 33.97 37.79
Fairness evaluation
According to the definition of fairness in FL from (Li et al. 2021), we can take the std of test accuracy as a fairness measure. Here we use the std of testing accuracy across 100 clients with the Cifar10 dataset as shown in the following table. Besides, we compare the local training time (in seconds) of each client in one round in the table below.
Metric FedAvg HeteroFL SplitMix FeDepth $m$-FeDepth
Time (s) 0.42 $\pm$ 0.05 0.75 $\pm$ 0.09 1.90 $\pm$ 0.60 2.49 $\pm$ 0.93 2.32 $\pm$ 0.93
Fairness 0.05253 0.03888 0.04919 0.04596 0.04762